Suppose that a hair salon owner in a small town has three workers. Suppose that one client


requires the whole day to be attended due to the complexity of the styles offered by the salon


owner. Suppose that the demand is 3 requests for hair to be done per day.


(a) Find the probability that the workers are idle in a given day.


(b) The probability that the salon owner has at least one client in a day.


(b) The probability that the salon owner has excess clients in a given day.

1
Expert's answer
2021-11-12T15:57:25-0500

Poisson distribution


"P(x=k)=\frac{\lambda^ke^{-\lambda}}{k!}"


mean:

"\lambda=3" clients per day


a)

"P(x=0)=e^{-3}=0.0498"


b)

"P(x\ge1)=1-P(x=0)=1-1.0498=0.9502"


c)

"P(x>3)=1-P(x=0)-P(x=1)-P(x=2)-P(x=3)"

"P(x=1)=3e^{-3}=0.1494"

"P(x=2)=3^2e^{-3}\/2=0.2241"

"P(x=3)=3^3e^{-3}\/6=0.2241"

"P(x>3)=1-0.0498-0.2241-0.2241=0.5020"